Joe invested $25,000 in an account that will draw 6.9% interest compounded quarterly for the next 10 years while Susan invests $25,000 in an account that will draw 6.8% interest compounded continuously over the next 10 years. Who will have the most money at the end of the 10 years?

What is the effective interest rate for 8% interest compounded daily?
Answer #1
Joe Investment future value= 25000*(1+6.9%/4)^(4*10)= $ 49,550,49
Susan Investment future value= 25000*(1+6.8%/365)^(365*10)=$ 49,343.82 (assumed daily compounding)
Joe will have more amount After 10 years
Effective interest rate=(1+8%/365)^365-1= 8.33%